OpenStudy (anonymous):

Identify the axis of symmetry for the graph of f(x) = (x - 3)2 + 1.

OpenStudy (precal):

x=3

hero (hero):

There's two ways two find axis of symmetry, although, there may be more... 1. Insert x = 3 to make the binomial (x-3)^2 = 0, then you know that the vertex will be (3,1) 2. Square the binomial and add 1 to get x^2 + 6x + 9 + 1 = x^2 + 6x + 10. Then use x = -b/2a to find the x axis of symmetry. You will get x = 3